OpenStudy (anonymous):

An 80.0 g arrow, initially at rest, is tiered from a bow whose string ecertain an average force of 95 N into the arrow over a distance of 80.0 cm. What is the speed of the arrow as it leaves the bow

OpenStudy (anonymous):

Work = force x distance W = F d = (95N)(0.80m) W = increase in kinetic energy = (1/2) m v^2 F d = (1/2) m v^2 substitute and solve.